OpenStudy (anonymous):

Which equation of a line has a slope of 3 and a y-intercept of -2? A. y = 3x - 2 B. y = 3x + 2 C. y = -3x + 2 D. y = -2x + 3

pooja195 (pooja195):

y=mx+b

OpenStudy (anonymous):

y = mx + b Where 'm' is the slope and 'b' is the y-intercept.

OpenStudy (anonymous):

So plug in '3' for 'm' and '-2' for 'b' and you have your equation.

OpenStudy (anonymous):

The equation is y = mx + b where m is slope and b is y-intercept, so can you figure out the answer by plugging in the values in the question? @lolcat

OpenStudy (anonymous):

so A

OpenStudy (anonymous):

Yep!

OpenStudy (luigi0210):

Also, y-intercept is when x=0 :P

OpenStudy (anonymous):

Yes, the answer is A! Great job! :)

OpenStudy (anonymous):

Or just where the line touches the y-axis. @Luigi0210

OpenStudy (anonymous):

Which equation of a line has a slope of -2 and a y-intercept of 4? A. y = -2x - 4 B. y = -2x + 4 C. y = 2x + 4 D. y = 4x - 2

OpenStudy (anonymous):

One question per post, close this one and open a new one.

OpenStudy (anonymous):

Yes, you can only ask 1 question on 1 post to prevent spam.